1.     With an 80% learning curve and an initial time of 200 minutes to do a job, at what number of repetitions will the job take less than 120 minutes?

Select one:

a. By the 6th necklace

b. By the 5th necklace

c. By the 7th necklace

d. By the 4th necklace
Answer #1

Learning rate equation is -

Tn=T1nb

where,
n = the unit number
T1 = time taken to produce first unit = 200 mins
Tn = time taken to produce nth unit
p = the learning rate = 0.80
b = In p / ln 2 = ln 0.80 / ln 2

Let the job n take less than 120 mins

=> Tn=T1nb

=> T1nb = 120

=> 200*n(ln 0.80 / ln 2 ) = 120

=> n(ln 0.80 / ln 2 ) = 120/200

=> n = (120/200)1/(ln 0.80 / ln 2 ) = 4.88

Hence, Time for 5th necklace will be lesser than 120 mins

Correct option (b) By the 5th necklace